OpenStudy (katherinesmith):

how do i simplify the square root of 404?

OpenStudy (katherinesmith):

like how do i break it down to a simpler square root

OpenStudy (tkhunny):

You must acquaint yourself with the idea of the "Prime Factorization". Ever hear of it?

OpenStudy (katherinesmith):

somewhat

OpenStudy (tkhunny):

Fair enough. Take 404 and find ANY factor. Name one.

OpenStudy (katherinesmith):

2

OpenStudy (tkhunny):

404 = 2*202 Is 2 prime? Is 202 prime?

OpenStudy (katherinesmith):

no, that can be divided by 2, and then you get 101

OpenStudy (katherinesmith):

which i think is prime

OpenStudy (tkhunny):

404 = 2*2*101 Are you SURE 101 is prime?

OpenStudy (katherinesmith):

yeah

OpenStudy (tkhunny):

The we're almost done: \(\sqrt{404} = \sqrt{2*2*101} = \sqrt{2^{2}\cdot 101} = \sqrt{2^{2}}\cdot \sqrt{101}\) One more step. Do you see it?

OpenStudy (katherinesmith):

the square root of 4 times 101

OpenStudy (katherinesmith):

so 2 to the square root of 101

OpenStudy (tkhunny):

Why did you go back to 4? We found the prime factors so we wouldn't have to say that any more. \(\sqrt{2^{2}}\cdot\sqrt{101} = 2\sqrt{101}\) No more 4. :-)

OpenStudy (tkhunny):

You beat me to it. Excellent work.

OpenStudy (tkhunny):

It is hoped that you do NOT think of 4 when simplifying \(\sqrt{2^{2}}\). It's just 2. No need to go through 4.

OpenStudy (katherinesmith):

okay. i hate this stuff. it's too much thinking.

OpenStudy (tkhunny):

It will be your friend, one day. It will save you at a car dealer or when your kids are in high school. One day, you will appreciate at least some small part. Good work and good luck.